5 ∫ 0 (x + 1) dx

50 (x + 1) dx

Grade:12

2 Answers

Miryala Gopalakrishna
9 years ago
integration of x+1  from 0 t0 5 is equal to
((x^2)/2)+x) from 0 to 5
=((5^2)/2+5)-0
=((25/2)+5)
=35/2
=17.5
Anusha
9 years ago
Ans:-50 (x+1)dx
=[(x2 /2)+x]05
=(25/2)+5
=(25+10)/2
=35/2